OpenStudy (anonymous):

Sabrina has $3.10 in nickels and dimes. If she has 52 coins in all, how many of each does he have?

OpenStudy (radar):

Here are the facts: 52 coins consisting of dimes and nickels. They add up to $3.10 Here is a way to solve a problem like this. Let x equal the number of dimes. (This is an arbitrary decision on my part) Since x equals the number of dimes how many nickels does she have? The total number of coins are 52. So the number of nickels are 52-x What do we have so far? x=number of dimes 52-x=number of nickels. Please note this is the number of dimes and nickels not their value. The value amount is important we know the value of 1 dime or nickel. simply multiply the value by the number of coins. this will get us the following

OpenStudy (radar):

.10 x is the value of dimes .05*(52-x) is the value of the nickels. Their combined total is $3.10 .10x + .05(52-x)=$3.10 ,10x +2.6-.05x=3.10 .05x=3.10-2.6=.5 or .05x=.5 number of dimes would be .5/.05=10 dimes. 52-10 gives the number of nickels, or 42 nickels. Lets check: Total number of coins is 10 + 42 = 52 coins. Check Now the amount of money, does it check? 10 dimes is $1.00, 42 nickels is $2.10, add them and the sum is $3.10 checks

OpenStudy (radar):

Did you follow? don't you just love word problems?!!!lol

OpenStudy (anonymous):

yeah thank you this really helped :)